OpenStudy (anonymous):

How many units apart are the points (4, -3) and (-4, 3)?

OpenStudy (anonymous):

Can someone please explain?

OpenStudy (debbieg):

Use the distance formula to find the distance between two points: \[d=\sqrt{(X_{1}-X_{2})^{2}+(Y_{1}-Y_{2})^{2}}\] Just plug in your point coordinates and simplify.

OpenStudy (anonymous):

so 4-(-4) ^+ (-3)-3^ ????????? then do the square root of what I get?

OpenStudy (debbieg):

You need to square each of those differences (I think that's what you mean, but it's hard to tell with just text).

OpenStudy (debbieg):

Tell me what you get, and I'll tell you if it's right. :)

OpenStudy (anonymous):

Would it be 64+9=73 then I square root 73?

OpenStudy (debbieg):

Check that 9.

OpenStudy (debbieg):

\((-3-3)^2=?\)

OpenStudy (anonymous):

um...0 ?

OpenStudy (debbieg):

What's -3-3?

OpenStudy (anonymous):

-3 - 3 = -6

OpenStudy (anonymous):

so 58 ?

OpenStudy (debbieg):

Good. now square that...? And add to 64

OpenStudy (debbieg):

No, look at the equation again. You squared the difference of x's but are forgetting to square the difference of the y's. :)

OpenStudy (anonymous):

oh so that means I have to square 64 then too right?

OpenStudy (debbieg):

No, the 64 is already the square!

OpenStudy (anonymous):

OH I GET WHAT U MEAN

OpenStudy (anonymous):

I was getting confused with the square root ,and squared

OpenStudy (debbieg):

Just take it one step at at time. :)

OpenStudy (anonymous):

3364 ?

OpenStudy (debbieg):

Oh lord. how did you get that? lol

OpenStudy (anonymous):

58x58 i dunno!

OpenStudy (anonymous):

what do I do!!?!?

OpenStudy (debbieg):

sorry, i'm having technical issues.... hang on.

OpenStudy (anonymous):

Oh okay (:

OpenStudy (debbieg):

\[d=\sqrt{(-4-4)^{2}+(3-(-3))^2}\] \[=\sqrt{(-8)^{2}+(6)^2}\] You basically got this far (notice I changed the order, just to be consistent with the formula in terms of calling the the point "1" and the other point "2"... but the order you subtract the coordinates in won't really matter, since it only changes the sign and we are going to square the result).

OpenStudy (debbieg):

NOW do the squares FIRST - they are under the square root - so you need \[(-8)^{2} ~~\text{and}~~~ 6^2\]

OpenStudy (anonymous):

-64 ? and 36?

OpenStudy (debbieg):

No real number, squared, is negative

OpenStudy (debbieg):

(but the 36 is correct)

OpenStudy (anonymous):

so its just left as 64 and 36

OpenStudy (debbieg):

remember, it's \((-8)^2\), NOT \(-8^2\), there is an important difference!

OpenStudy (debbieg):

Well, yes, it's "left" as those because that's what they ARE! :) OK, so now you have: \[\Large =\sqrt{64+36}\] Do the sum. THEN take the root.

OpenStudy (anonymous):

64+36=100 then rooted=10

OpenStudy (debbieg):

Tah-dah! :)

OpenStudy (anonymous):

So thats my answer? 10 Units away?

OpenStudy (debbieg):

yeppers, that's what the distance formula tells you. :)

OpenStudy (anonymous):

wow thanks (: explained very well...I dunno how to give you more than 1 medal id give you like 5 or something if I could

OpenStudy (debbieg):

Haha, thanks, one is fine. :) Happy to have helped, good job sticking with it and working through the explanation! :)